OpenStudy (anonymous):

quadratic equations : solve : (x+1)^2 = 100?

OpenStudy (anonymous):

x + 1 = pm 10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\( x=9,-11 \)

OpenStudy (anonymous):

\[x + 1 = \pm 10\]

OpenStudy (anonymous):

x = 9, x = -11